OpenStudy (pphalke):

how many liters of a 70% acid question must be mixed with a 20% acid solution to get 350 L of a 60% acid solution? plz help

OpenStudy (pphalke):

first person to solve this will get a medal

OpenStudy (welshfella):

x = volume of 70% solution and y = volume of the 20% so we can write x + y = 350

OpenStudy (welshfella):

we can create another equation using amounts of pure acid 0.7x is the volume of pure acid in the 70% solution 0.70x + 0.20y = 0.60*350 0.7x + 0.20y = 210

OpenStudy (pphalke):

how do i solve this?

OpenStudy (welshfella):

to solve for x and y the first thing you do is multiply the second equation by -5: x + y = 350 -------(1) 0.70x + 0.20y = 210 Multiply whole equation by -5 -3.5x - y = -1050 -- (2) ow add equations (1) and (2) . This will eliminate y and you can find value of x

OpenStudy (pphalke):

how do i add those ?

OpenStudy (08surya):

(70%)x+(20%)(350-x)=350*60% solve this

OpenStudy (08surya):

(70%)x+(20%)(350-x)=350*60% 70x+ 7000-20x=21000 50x=14000 5x=1400 x=280

OpenStudy (pphalke):

so @08surya ur saying that you need 280 liters to get 350 liters of a 60% acid solution

OpenStudy (pphalke):

sorry 70% acid solution

OpenStudy (08surya):

yup......m i wrong?

OpenStudy (welshfella):

280 L of 70% acid is correct